# Blowouts of Nascent Wind Bubbles in Pulsar-Driven Supernovae

### v1

@@ III.1.1 A Fiducial SESNe Blowout
Motivated by observations that such SNe can be well described by blackbody emission, at least up to and around the main peak (e.g., Leloudas et al., 2012; Nicholl and Smartt, 2016; Moriya et al., 2018), we use the bolometric luminosity and graybody temperature derived above to compute synthetic light curves in realistic survey bandpasses. In our fiducial SESN, the photospheric temperature remains above $\sim 10^{4}\,$K for an extended period and reaches $\gtrsim 10^{5}\,$K near the first peak, indicating that pulsar-driven SNe with wind-bubble blowout could, in some cases, produce detectable ultraviolet and soft X-ray emission. For additional illustrative cases and band-specific predictions, we refer to Section III.1.3.
@@ III.1.2 LSQ14bdq
Below we apply our blowout model to multi-band light curves of LSQ14bdq. In addition to the La Silla QUEST photometry converted to an SDSS $r$-like filter, we include the nearly contemporaneous $g$, $i$, and $z$-band measurements reported by Nicholl (2021) in the fit. The resulting best-fit model is shown in Fig. 5, obtained for $M_{\rm ej}\simeq 20\,M_{\odot}$, $E_{\rm sn}\simeq 2\times 10^{51}\,{\rm erg}$, $B_{\rm dip}\simeq 4.35\times 10^{13}\,{\rm G}$, and $P_{\rm NS}\simeq 1.56\,{\rm ms}$. These parameters are broadly consistent with the data. The model first peak is modestly faster and brighter than observed, plausibly due to the $\sim$2-day cadence (Nicholl et al., 2015b) that could have missed a brief, brighter maximum, to proximity to the marginal blowout boundary (Section IV.2).
@@ III.1.3 Diverse Blowout Signatures in SESNe
We now examine how variations in the pulsar engine parameters imprint diverse signatures on SESNe, fixing the ejecta mass at $M_{\rm ej}=10\,M_{\odot}$ and the explosion energy at $E_{\rm sn}=10^{51}\,{\rm erg}$. To provide a more representative view beyond a single case, Fig. 6 presents survey-ready light curves for four corner points of the blowout-enabled parameter space with the considering SESNe condition. These models span the combinations of high/low spin-down power and high/low rotational energy: (i) high $L_{\rm sp}$ and high $E_{\rm rot}$ ($B_{\rm dip}=10^{15}\,{\rm G}$, $P_{\rm NS}=1\,{\rm ms}$), (ii) low $L_{\rm sp}$ and high $E_{\rm rot}$ ($B_{\rm dip}=2\times 10^{13}\,{\rm G}$, $P_{\rm NS}=1\,{\rm ms}$), (iii) high $L_{\rm sp}$ and low $E_{\rm rot}$ ($B_{\rm dip}=10^{15}\,{\rm G}$, $P_{\rm NS}=3\,{\rm ms}$), and (iv) low $L_{\rm sp}$ and low $E_{\rm rot}$ ($B_{\rm dip}=2\times 10^{14}\,{\rm G}$, $P_{\rm NS}=3\,{\rm ms}$). As shown in the figures, the optical (ZTF) and ultraviolet (UVEX) bands display clear double-peaked morphologies in most samples, except in the UV for the $B_{\rm dip}=10^{15},$G, $P_{\rm NS}=3,$ms case; as we will show below in Fig. 9, this high-$B_{\rm dip}$, short-$P_{\rm NS}$ engine is more hypernova-like, with a cooler main peak that can suppress the second UV peak. In most cases, although the bolometric luminosity at the first peak is very high, the photospheric temperature is also high, placing the optical bands on the Rayleigh–Jeans tail of the spectral energy distribution (SED). Consequently, the first peak is typically far less prominent in the optical than in the UV. By contrast, models with lower $L_{\rm sp}$ are expected to have lower post-blowout graybody temperatures and therefore can display a more distinct optical double peak. In all cases, the early first peak evolves rapidly, so that at larger distances it is likely to be missed in purely optical searches. This in turn suggests that the intrinsic fraction of blowout events with double-peaked light curves among SLSNe-I may be higher than currently inferred from optical samples alone. Future UV-capable facilities such as UVEX could substantially increase the detection rate of such events and provide much richer diagnostic information.
@@ IV.2 Marginal Blowout
In Section III.1.2, our fit yields a first peak that is slightly faster and brighter than inferred from the data. The present framework also struggles to reproduce events with markedly slower first peaks. We do not view this as a failure of the blowout picture, but rather as a limitation of our current initialization of the pre-blowout phase. We assume that the swept-up shell evolves nearly adiabatically prior to blowout, with only minor energy losses through shock heating (e.g., Chevalier and Fransson, 1992). This approximation works best for strongly driven cases (high $B_{\rm dip}$ and short $P_{\rm NS}$), but it can bias solutions in the marginal-blowout regime, where radiative leakage may already be non-negligible before $t_{\rm bo}$ (Eq. 39).
Besides, to keep the adopted initial conditions self-consistent, we impose an additional admissibility cut: at $t_{\rm bo}$ we require the diffusion radius to lie outside the outer boundary, $R_{\rm diff}\geq R_{\rm out}$. This criterion is stricter than the physical blowout requirement. It excludes realizations that lie within the nominal blowout-allowed region of Fig. 2 but have $R_{\rm diff}<R_{\rm out}$ at $t_{\rm bo}$, which would imply appreciable radiative effects already during the pre-blowout evolution. Physically, however, blowout may still proceed even if the diffusion front penetrates the admixture layer before $t_{\rm bo}$, provided the radiative losses remain modest compared with the deposited energy.
@@
IV.3 Nonthermal Emission

### v2

@@ III.1.1 A Fiducial SESNe Blowout
Motivated by observations that such SNe can be well described by blackbody emission, at least up to and around the main peak (e.g., Leloudas et al., 2012; Nicholl and Smartt, 2016; Moriya et al., 2018), we use the bolometric luminosity and graybody temperature derived above to compute synthetic light curves in realistic survey bandpasses. In our fiducial SESN, the photospheric temperature remains above $\sim 10^{4}\,$K for an extended period and reaches $\gtrsim 10^{5}\,$K near the first peak, indicating that pulsar-driven SNe with wind-bubble blowout could, in some cases, produce detectable UV and soft X-ray emission. For additional illustrative cases and band-specific predictions, we refer to Section III.1.3.
@@ III.1.2 LSQ14bdq
Below we apply our blowout model to multi-band light curves of LSQ14bdq. In addition to the La Silla QUEST photometry converted to an SDSS $r$-like filter, we include the nearly contemporaneous $g$, $i$, and $z$-band measurements reported by Nicholl (2021) in the fit.
In the blowout framework, the two peaks are not fitted independently but are coupled by a small set of physical constraints that largely break parameter degeneracies. The main peak fixes the ejecta scale through $M_{\rm ej}$ and the total energy available up to peak, $E_{\rm tot}\simeq E_{\rm sn}+E_{\rm rot}$ (Kashiyama et al., 2016). The color evolution around the main peak provides an additional handle on the engine characteristics, because $E_{\rm rot}$ and $L_{\rm sp}$ imprint on $T_{\rm eff}(t)$ (e.g., Fig. 9; Kashiyama et al. (2016)). The first peak then constrains the early engine through the combination $E_{\rm sn}/L_{\rm sp}$ that sets its characteristic timescale and luminosity as described in Section II.3.4, while the solution must also lie in the blowout-enabled region of parameter space.
@@
The resulting best-fit model is shown in Fig. 5, obtained for
$M_{\rm ej}\simeq 20\,M_{\odot}$, $E_{\rm sn}\simeq 1\times 10^{51}\,{\rm erg}$, $B_{\rm dip}\simeq 3.69\times 10^{13}\,{\rm G}$, and $P_{\rm NS}\simeq 1.77\,{\rm ms}$. These parameters are broadly consistent with the data. The model first peak is modestly faster and brighter than observed, plausibly due to the $\sim$2-day cadence (Nicholl et al., 2015b) that could have missed a brief, brighter maximum, and because the inferred engine lies close to the marginal blowout boundary (Section IV.2), where the assumptions underlying our initial conditions become marginal while the same underlying physical picture still applies.
@@ III.1.3 Diverse Blowout Signatures in SESNe
We now examine how variations in the pulsar engine parameters imprint diverse signatures on SESNe, fixing the ejecta mass at $M_{\rm ej}=10\,M_{\odot}$ and the explosion energy at $E_{\rm sn}=10^{51}\,{\rm erg}$. To provide a more representative view beyond a single case, Fig. 6 presents survey-ready light curves for four corner points of the blowout-enabled parameter space with the considering SESNe condition. These models span the combinations of high/low spin-down power and high/low rotational energy: (i) high $L_{\rm sp}$ and high $E_{\rm rot}$ ($B_{\rm dip}=10^{15}\,{\rm G}$, $P_{\rm NS}=1\,{\rm ms}$), (ii) low $L_{\rm sp}$ and high $E_{\rm rot}$ ($B_{\rm dip}=2\times 10^{13}\,{\rm G}$, $P_{\rm NS}=1\,{\rm ms}$), (iii) high $L_{\rm sp}$ and low $E_{\rm rot}$ ($B_{\rm dip}=10^{15}\,{\rm G}$, $P_{\rm NS}=3\,{\rm ms}$), and (iv) low $L_{\rm sp}$ and low $E_{\rm rot}$ ($B_{\rm dip}=2\times 10^{14}\,{\rm G}$, $P_{\rm NS}=3\,{\rm ms}$). As shown in the figures, the optical (ZTF) and UV (UVEX) bands display clear double-peaked morphologies in most samples, except in the UV for the $B_{\rm dip}=10^{15},$G, $P_{\rm NS}=3,$ms case; as we will show below in Fig. 9, this high-$B_{\rm dip}$, short-$P_{\rm NS}$ engine is more hypernova-like, with a cooler main peak that can suppress the second UV peak. In most cases, although the bolometric luminosity at the first peak is very high, the photospheric temperature is also high, placing the optical bands on the Rayleigh–Jeans tail of the spectral energy distribution (SED). Consequently, the first peak is typically far less prominent in the optical than in the UV. By contrast, models with lower $L_{\rm sp}$ are expected to have lower post-blowout graybody temperatures and therefore can display a more distinct optical double peak. In all cases, the early first peak evolves rapidly, so that at larger distances it is likely to be missed in purely optical searches. This in turn suggests that the intrinsic fraction of blowout events with double-peaked light curves among SLSNe-I may be higher than currently inferred from optical samples alone. Future UV-capable facilities such as UVEX could substantially increase the detection rate of such events and provide much richer diagnostic information.
@@ IV.2 Marginal Blowout
In Section III.1.2, our fit yields a first peak that is slightly faster and brighter than inferred from the data. The present framework also struggles to reproduce events with markedly slower first peaks. We do not view this as a failure of the blowout picture, but rather as a limitation of our current initialization of the pre-blowout phase. We assume that the swept-up shell evolves nearly adiabatically prior to blowout, with only minor energy losses through shock heating (e.g., Chevalier and Fransson, 1992). This approximation works best for strongly driven cases (high $B_{\rm dip}$ and short $P_{\rm NS}$), but it can bias solutions in the marginal-blowout regime, where radiative leakage may already be non-negligible before $t_{\rm bo}$ (Eq. 39). This is related to the “inner-breakout” regime discussed by Kasen et al. (2016), in which diffusion begins to affect the dynamics and the emerging emission while the swept-up shell is still within the inner ejecta.
More specifically, to keep the adopted initial conditions self-consistent, we impose an additional admissibility cut: at $t_{\rm bo}$ we require the diffusion radius to lie outside the outer boundary, $R_{\rm diff}\geq R_{\rm out}$. This criterion is stricter than the physical blowout requirement. It excludes realizations that lie within the nominal blowout-allowed region of Fig. 2 but have $R_{\rm diff}<R_{\rm out}$ at $t_{\rm bo}$, which would imply appreciable radiative effects already during the pre-blowout evolution. Physically, however, blowout may still proceed even if the diffusion front penetrates the admixture layer before $t_{\rm bo}$, provided the radiative losses remain modest compared with the deposited energy.
@@
IV.3 56Ni Decay
We have neglected heating from 56Ni decay so far, since the central engine injects an energy comparable to the initial explosion energy by the onset of blowout, so radioactive input is typically subdominant. While, 56Ni is a key power source in supernova light curves, and its potential impact must be assessed.
For the first peak, 56Ni is expected to be synthesized predominantly in the deepest layers of the ejecta (Janka, 2012; Imasheva et al., 2023). Radioactive heating prior to blowout can, in principle, add energy to the swept up shell from smaller radii and reduce the energy injection required from the central engine. Nonetheless, this effect is negligible in practice. For a pure 56Ni, the total decay energy is $\epsilon_{\rm nuc}\simeq 0.11$ MeV/baryon. Considering the realistic 56Ni yields with $X_{\rm Ni}\lesssim\mathcal{O}(0.1)$ and a blowout timescale shorter than the relevant decay timescales, the radioactive energy that can be deposited before $t_{\rm bo}$ is limited to $\lesssim\mathcal{O}(0.01)$ MeV/baryon. This is below the characteristic energy scale required to accelerate the shell and trigger blowout, which is of order the ejecta specific kinetic energy, $\epsilon_{\rm kin}\sim 0.1$–0.5 MeV/baryon, for expansion velocities $v_{\rm ej}$ of $\sim 3{,}000$$10{,}000\ {\rm km\ s^{-1}}$. The radiation released around the first peak is supplied mainly by two channels: internal energy injected and stored prior to blowout and shock heating during blowout. Both operate at a scale $\sim\epsilon_{\rm kin}$. Therefore, within the blowout branch considered here, 56Ni heating does not appreciably modify the shell dynamics, the blowout time $t_{\rm bo}$, or the luminosity and timescale of the first peak. The same conclusion applies to SESNe and to lower-energy explosions, including ultrastripped and AIC/MIC SNe.
The role of 56Ni during the main peak is more parameter dependent and can be evaluated with simple scalings. Kashiyama et al. (2016) estimate the peak luminosity powered by 56Ni decay as
$L_{\rm sn}^{\rm Ni}\sim 10^{42}\ {\rm erg\ s^{-1}}\left(\frac{M_{{}^{56}{\rm Ni}}}{0.1\,M_{\odot}}\right)\left(\frac{M_{\rm ej}}{10\,M_{\odot}}\right)^{-1}\left(\frac{v_{\rm ej}}{10^{9}\ {\rm cm\ s^{-1}}}\right)\left(\frac{\kappa}{0.1\ {\rm cm^{2}\ g^{-1}}}\right)^{-1},$
and the peak luminosity in the pulsar driven case as
$L_{\rm sn}^{\rm psr}\sim 3\times 10^{44}\ {\rm erg\ s^{-1}}\left(\frac{B_{\rm dip}}{10^{14}\ {\rm G}}\right)^{-2}\left(\frac{M_{\rm ej}}{10\,M_{\odot}}\right)^{-1}\left(\frac{v_{\rm ej}}{10^{9}\ {\rm cm\ s^{-1}}}\right)\left(\frac{\kappa}{0.1\ {\rm cm^{2}\ g^{-1}}}\right)^{-1}.$
For the SESN scalings adopted in this paper, these estimates give $L_{\rm sn}^{\rm psr}\gtrsim 10\,L_{\rm sn}^{\rm Ni}$ when $B_{\rm dip}\gtrsim 3\times 10^{13}\ {\rm G}$, so the radioactive contribution to the peak luminosity is limited. At lower $B_{\rm dip}$, the relative contribution from 56Ni increases and overlaps with the low $L_{\rm sp}$ parameter space associated with marginal blowout. For low mass, low energy AIC or MIC explosions, a similar estimate indicates that $B_{\rm dip}\gtrsim$ a few $\times 10^{12}\ {\rm G}$ often still gives $L_{\rm sn}^{\rm psr}\gtrsim 10\,L_{\rm sn}^{\rm Ni}$, whereas for weaker engines near marginal blowout the radioactive contribution to the main peak can become non negligible. Since we explicitly exclude marginal blowout in Section IV.2, radioactive heating does not affect the calculations presented here. A self consistent inclusion of 56Ni decay is required when extending the model to the marginal blowout regime.
After the main peak, as the ejecta becomes progressively more transparent, $\gamma$-ray escape from the 56Ni decay chain increases and the deposition efficiency declines. Positron deposition, mainly from the $\beta^{+}$ branch of 56Co decay, can then become relatively more important, which changes the late time deposition channel and can modify the optical output at specific phases (Milne et al., 1999, 2001; Seitenzahl, 2011). The formation of an admixture layer may also enhance outward mixing and transport some 56Ni to larger radii, which would promote earlier $\gamma$-ray leakage and alter the late time decline, potentially introducing non thermal signatures. In summary, within our present framework that focuses on thermal emission and excludes marginal blowout, neglecting 56Ni heating is an acceptable approximation. Incorporating 56Ni decay self consistently is a natural next step for extensions that include marginal blowout and late time emission.
IV.4 Nonthermal Emission
@@
IV.5 Comparison with alternative early-peak mechanisms
Several alternatives to blowout can also produce a double-peaked morphology, including shock cooling of extended material and CSM interaction in some configurations (e.g., Nakar and Piro, 2014; Chevalier and Irwin, 2011). To enable observational separation, we compare their predicted timescales, temperature, color evolution, and multiband behavior against those expected from a single ejecta and engine system.
First, blowout admits a stringent internal-consistency test because the first peak and the main peak originate from the same ejecta and engine system. A single parameter set must account for both peaks, with the bulk scale primarily set by $(M_{\rm ej},\,E_{\rm sn},\,E_{\rm rot},\,L_{\rm sp})$. By contrast, CSM-based explanations introduce additional degrees of freedom (e.g., CSM mass and radius) that are only weakly constrained and can vary widely across events (Piro, 2015; Fraser, 2020; Smith, 2014).
Second, the radiated-energy budget of the first peak is a robust discriminator. In shock-cooling models in which an extended low-mass layer reprocesses the explosion energy, the radiated energy is typically (Nakar and Piro, 2014; Piro, 2015)
$E_{\rm ext}\sim 9\times 10^{48}\ {\rm erg}\,\left(\frac{E_{\rm sn}}{10^{51}\ {\rm erg}}\right)\left(\frac{M_{\rm ej}}{10\ M_{\odot}}\right)^{-0.7}\left(\frac{M_{\rm ext}}{0.01\ M_{\odot}}\right)^{0.7},$ (67)
which, for the fiducial SESN parameters adopted in this paper, is naturally of order $10^{49}$ erg and can plausibly reach $\sim 10^{50}$ erg depending on the extended-material or CSM configuration (e.g., Nyholm et al., 2020; Hiramatsu et al., 2024). In blowout, the engine must deposit an energy comparable to the explosion energy before transparency and release it rapidly. The characteristic early energy scale therefore satisfies $E_{\rm bo}\gtrsim E_{\rm sn}$ (Section II.2), and is typically orders of magnitude larger than $E_{\rm ext}$ for comparable $M_{\rm ej}$ and $E_{\rm sn}$.
Third, for a fixed observed duration $t_{\rm first}$, this energy contrast maps directly into peak luminosity and hence multiband detectability. For shock cooling, a characteristic luminosity scale is $L_{\rm ext}\sim E_{\rm ext}/t_{\rm first}\sim 10^{44}\ {\rm erg\,s^{-1}}\,(E_{\rm ext}/10^{49}\ {\rm erg})\,(t_{\rm first}/{\rm day})^{-1}$. In blowout, the first-peak luminosity is set by the energy stored in the admixture layer and released over the broadening timescale, $L_{\rm first}\sim E_{\rm kin,adm}/t_{\rm bro}$, and asymptotes to the limiting regimes in Eqs. (63)–(64). For fiducial SESN parameters this corresponds to $L_{\rm first}\sim 10^{46}\ {\rm erg\,s^{-1}}$, indicating that blowout can be brighter than shock cooling by orders of magnitude at the same $t_{\rm first}$.
Finally, the temperature and color evolution provide an additional discriminator. For shock cooling, the effective temperature depends sensitively on the radius of the extended material and can therefore span a broad range. For the same observed first-peak duration $t_{\rm first}$ and a comparable opacity normalization, a commonly used scaling is (Nakar and Piro, 2014)
$T_{\rm eff}\sim 4\times 10^{4}\ {\rm K}\,\left(\frac{\kappa}{0.1\ {\rm cm^{2}\,g^{-1}}}\right)^{0.25}\left(\frac{t_{\rm first}}{1\ {\rm day}}\right)^{-0.5}\left(\frac{R_{\rm ext}}{10^{13}\ {\rm cm}}\right)^{0.25}.$ (68)
Accordingly, CSM-involved scenarios do not necessarily guarantee a strong UV excess, and the first-peak optical colors can show substantial event-to-event diversity, largely reflecting the spread in $R_{\rm ext}$ and other CSM properties. In the blowout picture, the rapid early expansion largely suppresses sensitivity to the progenitor radius and tends to keep the first peak hot and blue. Evaluated under the same fiducial SESN parameter set and at the same $t_{\rm first}$, our analytic estimate gives $T_{\rm eff,1}\sim 10^{5}$ K (Eq. (65)), implying a prominent UV contribution and very blue optical colors.
Taken together, blowout is distinguished not merely by a double-peaked morphology, but by the combination of (i) a coupled two-peak fit under a single physically allowed parameter set, (ii) a large early energy scale set by $E_{\rm bo}\gtrsim E_{\rm sn}$, (iii) a brighter first peak, and (iv) a typically hot and UV-bright early color temperature. Early-time UV coverage and well-sampled multiband color evolution therefore provide the most direct observational tests. We emphasize that the diversity of CSM and extended-envelope configurations can lead to partial overlap in phenomenology for some parameter choices, and the separation is more robust when the two peaks provide a test of parameter consistency, with the energy scale, first-peak brightness, and color evolution jointly constrained.